what is the equation of the line through points (-3,2) and (-4,5)
Answer by CharlesG2(834)

what is the equation of the line through points (-3,2) and (-4,5)
slope m = rise/run = (y2 - y1)/(x2 - x1)
m = (5 - 2)/(-4 - -3)
m = 3/(-4 + 3)
m = 3/(-1)
m = -3
slope intercept form of a line is y = mx + b,
where b is the y-intercept or vertical-intercept or the point (0,b)
y = -3x + b
plug in either of the points to solve for b
plugging in (-3,2):
2 = -3(-3) + b
2 = 9 + b
-7 = b
plugging in (-4,5):
5 = -3(-4) + b
5 = 12 + b
again -7 = b
equation of the line:
y = -3x - 7